# Disentangling Learning from Judgment: Representation Learning for Open Response Analytics

### v1

@@ 1. Introduction
We propose a complementary, analytics-first approach. Rather than accepting automated scores as final, systems should render model behavior auditable by linking predictions to each teacher’s grading history and revealing where content signals diverge from rater priors. This is critical since teachers may accept the automated grades without much scrutiny (Xavier et al., 2025). Our contributions are twofold: (a) a deconfounded representation-learning framework that disentangles student content from teacher priors, and (b) early insights into practitioner-facing analytics that expose disagreement patterns, transforming embeddings from passive features into tools for inspecting open-response grading.
@@ 5. Discussion
We reframe automated analysis of open responses from a focus on predicting grades to a focus on making the grading process itself intelligible. Our central novelty claim is to separate content signals encoded in response embeddings from rater tendencies captured by teacher priors, then to read their points of convergence and divergence as evidence about what is being modeled. In doing so, we treat open responses not only as inputs for classification but also as pedagogically meaningful artifacts whose interpretation must remain visible to teachers and researchers (Gibson and Shibani, 2022; Botelho et al., 2023; Gurung et al., 2022).
@@
Addressing RQ2, accounting for teacher leniency and severity, sharpened the content signal rather than diluting it. After partialling out rater baselines and problem-level regularities through centering and residualization (Enders and Tofighi, 2007; Zhang et al., 2025), more embedding coordinates survived Lasso regularization, and discrimination improved at comparable or lower error. This pattern suggests that priors otherwise absorb variance that overlaps with content features; once that variance is allocated to explicit rater terms, the remaining coefficients concentrate on semantically coherent directions. The effect is methodological and interpretive. Methodologically, the adjusted design yields models that generalize better across teachers because the confounding pathway from rater practice to content proxies is attenuated. Interpretively, the nonzero embedding dimensions are easier to audit because they align with content that remains predictive even when rater baselines are held fixed. For learning analytics, these sharpened signals create opportunities to surface the linguistic and conceptual features of student work that most influence grading. Presenting such dimensions back to teachers can foster reflection and refinement of scoring practices. In future work, we plan to develop teacher-facing analytics that highlight these features to support calibration and alignment with evidence of student understanding.
@@ 5.1. Implications and Future Directions for Learning Analytics
Methodologically, this work demonstrates that linear models paired with principled residualization provide a transparent baseline for open-response analytics. Prior work has demonstrated the utility of embeddings for automated scoring (Erickson et al., 2020; Botelho et al., 2023); here, we extend this by showing that explicitly allocating variance to rater effects enables the residual content signal to generalize across teachers and produce interpretable artifacts for qualitative inspection.
@@ 6. Conclusion
We recast open-response analytics as the task of making judgment visible. Modeling teacher priors alongside response embeddings reveals that much of the predictive signal originates from the rater context; once this context is explicit, the residual content representation becomes sharper and more interpretable. Disagreement patterns between content-only, prior-only, and integrated models then serve as diagnostic evidence about what is being rewarded, allowing systems to surface cases where semantic understanding conflicts with historical grading tendencies.
Our contribution is a novel pipeline for creating learning analytics that can inform teacher- and researcher-facing tools in future work. This pipeline involves temporal validation with explicit rater baselines, orthogonalized embeddings, and a projection workflow for qualitative inspection, which transforms embeddings from mere features into tools for reflection. This allows the end-user to decompose predictions, enabling them to interrogate how student reasoning and their evaluation interact. These interrogations will contribute to making teacher- and AI-graded student open responses fairer and more aligned with cognitive strategies that are known to enhance learning, such as self-explanation.
Acknowledgments
The authors would like to thank the ASSISTments Foundation for the provision and curation of the QC23-Complete dataset. CB wishes to thank Kevin K. Tang for contributions to data preprocessing code as well as Vincent Aleven for co-supervision of MP. MP was supported by the Summer Undergraduate Research Apprenticeship (SURA) program at Carnegie Mellon University.

### v2

@@ 1. Introduction
We propose a complementary, analytics-first approach. Rather than accepting automated scores as final, systems should render model behavior auditable by linking predictions to each teacher’s grading history and revealing where content signals diverge from rater priors. This is critical because teachers may reasonably rely on automated grades for efficiency, which can reduce opportunities to inspect why a score was assigned (Xavier et al., 2025). Our contributions are twofold: (a) a deconfounded representation-learning framework that disentangles student content from teacher priors, and (b) early insights into practitioner-facing analytics that expose disagreement patterns, transforming embeddings from passive features into tools for inspecting open-response grading.
@@ 5. Discussion
We reframe automated analysis of open responses from a focus on predicting grades to a focus on making the grading process itself intelligible. Our central novelty claim is to separate content signals encoded in response embeddings from rater tendencies captured by teacher priors, then to read their points of convergence and divergence as evidence about what is being modeled. In doing so, we treat open responses as pedagogically meaningful artifacts whose interpretation must remain visible to teachers and researchers (Gibson and Shibani, 2022; Botelho et al., 2023; Gurung et al., 2022; Xavier et al., 2025).
@@
Addressing RQ2, accounting for teacher leniency and severity, sharpened the content signal rather than diluting it. After partialling out rater baselines and problem-level regularities (Enders and Tofighi, 2007; Zhang et al., 2025), more embedding coordinates survived Lasso regularization, and discrimination improved at comparable or lower error. This pattern suggests that priors otherwise absorb variance that overlaps with content features; once that variance is allocated to explicit rater terms, the remaining coefficients concentrate on semantically coherent directions. The effect is methodological and interpretive. Methodologically, the adjusted design yields models that generalize better across teachers because the confounding pathway from rater practice to content proxies is attenuated. Interpretively, the nonzero embedding dimensions are easier to audit because they align with content that remains predictive even when rater baselines are held fixed. For learning analytics, these sharpened signals create opportunities to surface the linguistic and conceptual features of student work that most influence grading. Presenting such dimensions back to teachers can foster reflection and refinement of scoring practices. In future work, we plan to develop teacher-facing analytics that highlight these features to support calibration and alignment with evidence of student understanding.
@@ 5.1. Implications and Future Directions for Learning Analytics
Methodologically, this work demonstrates that linear models paired with explicit rater covariates and simple embedding transformations provide a transparent baseline for open-response analytics. Prior work has demonstrated the utility of embeddings for automated scoring (Erickson et al., 2020; Botelho et al., 2023); here, we extend this by showing that explicitly modeling rater effects alongside content embeddings enables the content signal to generalize across teachers and produce interpretable artifacts for qualitative inspection.
@@ 6. Conclusion
We recast open-response analytics as the task of making judgment visible. Modeling teacher priors alongside response embeddings reveals that much of the predictive signal originates from the rater context; once this context is explicit, the remaining contribution of the response-based representations becomes easier to interpret and audit. Disagreement patterns between content-only, prior-only, and integrated models serve as diagnostic evidence about why a score was produced, allowing systems to surface cases where the content of a student’s response supports understanding but the predicted score is driven primarily by historical grading patterns. Such cases are natural candidates for teacher review, moderation, or discussion.
Our contribution is a novel pipeline for creating learning analytics that can inform teacher- and researcher-facing tools in future work. This pipeline involves temporal validation with explicit rater baselines and methods for combining these baselines with representation learning for qualitative inspection, which transforms embeddings into tools for reflection. This allows the end-user to decompose predictions, enabling them to interrogate how student reasoning and their evaluation interact. These interrogations will contribute to making teacher- and AI-graded student open responses fairer and more aligned with cognitive strategies that are known to enhance learning, such as self-explanation.
Acknowledgements.
The authors would like to thank the ASSISTments Foundation for the provision and curation of the QC23-Complete dataset. CB wishes to thank Kevin K. Tang for contributions to the data preprocessing code, as well as Vincent Aleven for co-supervision of MP. MP was supported by the Summer Undergraduate Research Apprenticeship (SURA) program at Carnegie Mellon University.